OpenStudy (anonymous):

help with this differential equation \frac{ dy }{ dt } = y^{2} - 2ty + t^{2} +1

OpenStudy (anonymous):

\[ \frac{ dy }{ dt } = y^{2} - 2ty + t^{2} +1\]

OpenStudy (anonymous):

\[\frac{ dy }{ dt } = y^{2} - 2ty + t^{2} +1\] You want to integrate?

OpenStudy (anonymous):

I want y

OpenStudy (anonymous):

I managed to do it , it was a simple substitution x = y-t and solve for x

OpenStudy (anonymous):

Excellent. Well Done.